A system consisted of 50 mg of an organic compound dissolved in 1.00 mL of water (solvent A). Compare the effectiveness of one 2.00 mL extraction versus two 1.00 mL extractions with ether (solvent B) where the distribution coefficient, K, is equal to 2. The answer gives you a fraction of the amount of organic compound extracted into solvent B, ether. How much material (in mg) is left in the water layer? Discuss which extraction method is preferable; one or two extractions?
1
Expert's answer
2013-03-13T07:39:27-0400
In this case distribution coefficient, K is concentration of compound in top phase/concentration of compound in bottom phase. Ratio of the volume is equal to the allocation between phases.

In first case V1:V2 = 2 : 1
Ratio of concentration of the organic compound in phases 2 : 1(water)
50/(2+2+1) *4 = 40 g of organic compound could be obtained from organic solvent

In second case first step V1:V2 = 1 : 1
Ratio of concentration of the organic compound in phases 2 : 1(water)
50/3 *2 = 33.33 g of organic compound could be obtained from organic solvent.

In second case second step V1:V2 = 1 : 1
Ratio of concentration of the organic compound in phases 2 : 1(water)
(50 - 33.33)/3 *2 = 11.11 g of organic compound could be obtained from organic solvent.

Total mass of organic compound is 44.44 g, so this method is more preferable.